# Multiple Imputation of Hierarchical Nonlinear Time Series Data with an Application to School Enrollment Data

### v1

@@ Abstract
International comparisons of hierarchical time series data sets based on survey data, such as annual country-level estimates of school enrollment rates, can suffer from large amounts of missing data due to differing coverage of surveys across countries and across times. A popular approach to handling missing data in these settings is through multiple imputation, which can be especially effective when there is an auxiliary variable that is strongly predictive of and has a smaller amount of missing data than the variable of interest. However, standard methods for multiple imputation of hierarchical time series data can perform poorly when the auxiliary variable and the variable of interest are have a nonlinear relationship. Performance of standard multiple imputation methods can also suffer if the substantive analysis model of interest is uncongenial to the imputation model, which can be a common occurrence for social science data if the imputation phase is conducted independently of the analysis phase. We propose a Bayesian method for multiple imputation of hierarchical nonlinear time series data that uses a sequential decomposition of the joint distribution and incorporates smoothing splines to account for nonlinear relationships between variables. We compare the proposed method with existing multiple imputation methods through a simulation study and an application to secondary school enrollment data. We find that the proposed method can lead to substantial performance increases for estimation of parameters in uncongenial analysis models and for prediction of individual missing values.
@@ 1 Introduction
Missing values within hierarchical time series data sets are a common occurrence in social science data, particularly for comparisons across countries and across times that rely on survey data. International surveys may only be conducted in selected years and coverage of countries may differ from year to year, while country-level surveys and censuses may be conducted annually in some countries but may only occur sporadically in others. The presence of missing data can be compounded in settings that require the compilation of data from multiple different sources. One example where this occurs is for school enrollment data, where the UNESCO Institute for Statistics compiles survey and administrative data to create annual, internationally comparable estimates of school enrollment rates. The differing availability of survey and administrative data for each country and year leads to a large number of country-years in the UNESCO database where enrollment data is missing. Researchers interested in questions comparing survey-based indicators like enrollment rates across countries and across times can thus encounter large amounts of missing hierarchical time series data.
For some hierarchical time series variables, there may be related variables that that measure similar underlying concepts but have different amounts of missing data. For example, this can occur if a survey is designed to first ask for more specific information, but, failing to obtain the specific information, then asks for more general information. If the variable of interest for analysis is the variable that has a greater amount of missing data, researchers might be interested in how to best leverage the information from the auxiliary variable that has less missing data to impute the variable of interest. This situation arises with school enrollment data, where two commonly reported measures of enrollment rates are the Net Enrollment Rate (NER) and the Gross Enrollment Ratio (GER). Measurement of NER requires knowledge of the number and age distribution for children who are enrolled, while measurement of GER only requires knowledge of the number of children enrolled. The comparative ease of measuring GER can result in more missing values for NER compared to GER. If researchers are more interested in analyses using NER, they may want to use the available information about the auxiliary variable GER to help impute missing values for NER using a multiple imputation procedure.
@@
Multiple imputation, first developed by Rubin (1978, 1987), is a widely used approach for handling missing data. In multiple imputation, $M>1$ imputed values for missing observations are sampled from the posterior predictive distribution of the missing data given the observed data. The $M$ imputed values result in $M$ completed data sets, each of which consists of the observed data and one set of imputed values for the missing observations. The completed data sets can each be analyzed separately using complete data methods and the results of the analyses can be combined into one final, pooled result using combining rules from Rubin (1987) that account for both within-imputation variation and between-imputation variation.
For hierarchical data, multiple imputation approaches that do not explicitly account for the hierarchical structure of the data can lead to biased results in downstream analyses (Taljaard et al., 2008; Enders et al., 2016; Lüdtke et al., 2017). Many approaches specifically for multiple imputation of hierarchical time series data have been developed (e.g. Liu et al., 2000; He et al., 2011; Speidel et al., 2018; Enders et al., 2020; Grund et al., 2021; among others), with two of the most widely used approaches for social science data being Amelia II and multilevel extensions of Multiple Imputation by Chained Equations (MICE). Amelia, originally developed by King et al. (2001) and extended as Amelia II by Honaker and King (2010), is a multiple imputation method designed specifically for hierarchical time series data. Amelia is based on the joint modeling approach to multiple imputation, where imputed values are sampled from a joint distribution for all variables with missing data. MICE is a multiple imputation method developed by van Buuren and Groothuis-Oudshoorn (2011) that uses the fully conditional specification (FCS) approach to multiple imputation. Rather than explicitly specifying a joint imputation model, the FCS algorithm iteratively samples from univariate conditional imputation models for all variables with missing data until convergence is reached. Several methods that account for hierarchical data structures have been implemented within the MICE framework, including the linear mixed effects method developed by Schafer and Yucel (2002). These commonly used methods assume that variables in the imputation model have a linear relationship. In settings where variables have a strong nonlinear relationship and transformation to approximate linearity is not possible, these methods are misspecified and can result in poor imputations.
@@
We focus on the setting where the imputation of missing values and the analysis of imputed values are conducted independently. The development of multiple imputation originated in this setting, where Rubin (1977, 1978) proposed the multiple imputation framework as a way to provide imputed values for missing responses in public-use releases of large data sets from sample surveys. An appealing feature of multiple imputation for practitioners is the ability to use the same imputed data set to conduct many different analyses (Schafer, 1997a; Rubin, 1987). However, using the same multiply imputed data for different analyses can lead to the analysis model being uncongenial to the imputation model in the sense of Meng (1994). The analysis model and the imputation model are congenial if the imputation model contains at least as much information as the analysis model. Many of the theoretical properties of multiply imputed estimates, such as the consistency of estimates using Rubin’s combining rules, rely on congeniality (Meng, 1994; Rubin, 1996; Xie and Meng, 2017). In practice, uncongeniality of the analysis and imputation models can be a regular occurrence when the researchers that collect the survey data create imputed values for publication that are then used in analyses by external researchers. The external researchers may not have access to the same information or resources needed to create imputations of their own, for example if the imputation process is not well-documented or if the predictive variables used in the imputation model are not publicly available. In this scenario, the analysis models used by the external researchers are not guaranteed to be congenial to the imputation model. The ability of a multiple imputation method to perform well for uncongenial analyses is thus of interest for practitioners.
In this paper, we propose a multiple imputation method for continuous hierarchical time series data that can account for nonlinear relationships between variables. We refer to this method as MINTS for Multiple Imputation of hierarchical Nonlinear Time Series data. We focus on the bivariate setting, where one variable is the variable of interest for which imputations are desired and the second variable is an auxiliary variable that is easier to measure and has a nonlinear relationship with the variable of interest. We also focus on a specific type of nonlinear relationship where the nonlinearity takes the form of a piecewise linear function. We evaluate the out-of-sample validation performance of MINTS using a simulation study that compares the performance of the proposed method with existing multiple imputation methods using simulated data, where we focus on estimating parameters in analysis models that are uncongenial to the imputation model. We also conduct two validation exercises using a motivating data set on secondary school enrollment rates, where we evaluate the predictive performance for estimating individual missing values and for estimating parameters in uncongenial analysis models. Finally, we report the imputation results using the full enrollment data and make available a data set including 40 multiple imputations for NER and GER.
@@
This paper is structured as follows. In Section 2, we describe the motivating case study of secondary school enrollment data in further detail. Section 3 describes the proposed multiple imputation method. In Sections 4 and 5, we evaluate how well the proposed multiple imputation method performs for estimation of parameters in uncongenial analysis models and for prediction of individual missing observations through a simulation study and an application to the enrollment data. Section 6 includes further discussion and comparison to existing imputation methods. Finally, we summarize the findings of this paper in Section 7.
@@ 2 Motivating Case Study: Secondary School Enrollment Rates
The UNESCO Institute for Statistics collects internationally comparable data on education indicators on an annual basis for all countries of the world, based largely on survey and administrative data (UNESCO Institute for Statistics, 2023). The World Bank combines this education data with population data from the United Nations Population Division to create estimates of two types of enrollment rates: the Net Enrollment Rate (NER) and the Gross Enrollment Ratio (GER) (World Bank, 2021). We focus on secondary school enrollment. NER is the ratio of children of official secondary school age who are enrolled in secondary school to the population of official secondary school age children. NER is bounded between 0% and 100% and both the numerator and denominator reflect children of official secondary school age. GER is the ratio of total enrollment in secondary school, regardless of age, to the population of official secondary school age children. The numerator and denominator for GER potentially represent different populations, where children who are not of official secondary school age can be counted in the numerator but not in the denominator. Thus, GER can be greater than 100% if children who are enrolled in secondary school are not of official school age. The two measures of enrollment are subject to the boundary NER $\leq$ GER and have a strong nonlinear relationship that can be well-approximated by a piecewise linear function.
For substantive analyses, one measure of enrollment may be preferred over the other. NER can be thought of as a demographic rate, where the numerator counts the number of enrollments for the population of children of official secondary school age in a given year and the denominator counts the person-years lived in that population for the given year. NER can thus be preferable over GER for demographic analyses. However, historical time series of NER tend to have more missing values than time series of GER. Measurement of NER is more difficult than measurement of GER due to requiring knowledge of the age distribution for children enrolled in secondary school. Measurement of GER is comparatively easy, as GER can be calculated using only the number of children who are currently enrolled. For school systems that do not have robust recordkeeping systems and countries that do not have good vital registration systems, knowledge of the age of all enrolled children can be difficult to obtain. The greater availability of estimates of GER and the strong relationship between NER and GER motivates the desire to impute missing values of NER using the relationship between NER and GER.
@@
Figure 1 shows a scatter plot of the complete cases for NER and GER. The red line illustrates the B-spline of degree 1 fit to the complete cases using the A-splines methodology described in Section 3.2.5. There is a nonlinear relationship between NER and GER, with a shift in trend occurring around GER $=100$. The variation of NER about the fitted spline also appears to vary with GER, with smaller variability at the lowest levels of GER and larger variability around GER $=100$.
Figure 1: Scatter plot of complete cases for NER and GER from secondary enrollment data set, with the B-spline of degree 1 fit using A-splines superimposed in red.
Time series of NER and GER are shown in Figure 2 for Afghanistan, Belgium, Spain, and Nigeria. Afghanistan is an example of the most common type of pattern seen for individual countries, where there is a larger number of observed values for GER compared to NER. Only one country, Brazil, has the opposite pattern with one more observed value of GER than observed values of NER. Belgium and Spain are two examples of countries where the nonlinear relationship between NER and GER is visible in the time series for the individual country. Both Belgium and Spain have relatively few missing values for GER, but have large stretches of time with no observations for NER. Finally, Nigeria is an example of a country that has some observed values for GER but has no observed values for NER. There are 14 countries in the enrollment data set that, like Nigeria, have at least one observation for GER but no observations for NER.
Figure 2: Observed values of NER and GER for selected countries from secondary enrollment data set. Solid black and red circles indicate observed values for GER and NER, respectively.
@@ 3.1 Notation
Let the matrices $\mathbf{R}^{X}$ and $\mathbf{R}^{Y}$ denote the response matrices for $\mathbf{X}$ and $\mathbf{Y}$, respectively. If an element $(c,t)$ is observed, then the corresponding element of the response matrix equals 1. For example, if $X_{c,t}$ is observed, then $R_{c,t}^{X}=1$. If $X_{c,t}$ is missing, then $R_{c,t}^{X}=0$. The matrices $\mathbf{X}$ and $\mathbf{Y}$ can also be written in terms of their observed and missing portions, e.g. $\mathbf{X}=[\mathbf{X}_{mis},\mathbf{X}_{obs}]$ where the observed portion of $\mathbf{X}$ is denoted $\mathbf{X}_{obs}$ and the unobserved portion of $\mathbf{X}$ is denoted $\mathbf{X}_{mis}$.
For the enrollment data, the auxiliary variable $\mathbf{X}$ is GER and the variable of interest $\mathbf{Y}$ is NER. The enrollment data includes a total of $C=202$ countries, where the assignment of countries to indices $c\in 1,...,C$ was done alphabetically by country name. There are $T=51$ years in the enrollment data set, where $t=1$ corresponds to the year 1970 and $t=51$ corresponds to the year 2020.

### v2

@@ Abstract
International comparisons of hierarchical time series data sets based on survey data, such as annual country-level estimates of school enrollment rates, can suffer from large amounts of missing data due to differing coverage of surveys across countries and across times. A popular approach to handling missing data in these settings is through multiple imputation, which can be especially effective when there is an auxiliary variable that is strongly predictive of and has a smaller amount of missing data than the variable of interest. However, standard methods for multiple imputation of hierarchical time series data can perform poorly when the auxiliary variable and the variable of interest have a nonlinear relationship. Performance can also suffer if the multiple imputations are used to estimate an analysis model that makes different assumptions about the data compared to the imputation model, leading to uncongeniality between analysis and imputation models. We propose a Bayesian method for multiple imputation of hierarchical nonlinear time series data that uses a sequential decomposition of the joint distribution and incorporates smoothing splines to account for nonlinear relationships between variables. We compare the proposed method with existing multiple imputation methods through a simulation study and an application to secondary school enrollment data. We find that the proposed method can lead to substantial performance increases for estimation of parameters in uncongenial analysis models and for prediction of individual missing values.
Multiple imputation, missing data, Bayesian hierarchical model, time series, school enrollment,
keywords:
[…]
, and
@@ 1 Introduction
Missing values within hierarchical time series data sets are a common occurrence in social science data, particularly for comparisons across countries and across times that rely on survey data. International surveys may only be conducted in selected years and coverage of countries may differ from year to year, while country-level surveys and censuses may be conducted annually in some countries but may only occur sporadically in others. The presence of missing data can be compounded in settings that require the compilation of data from multiple different sources. One example where this occurs is for school enrollment data, where the UNESCO Institute for Statistics compiles survey and administrative data from countries around the world to create annual, internationally comparable estimates of school enrollment rates. The country-specific time series for school enrollment rates have a hierarchical structure where countries are nested in the world. The differing availability of survey and administrative data across countries and years leads to a large number of country-years in the UNESCO database with missing data.
For hierarchical time series data arising from survey data, there can be multiple variables that measure similar underlying quantities but have differing amounts and patterns of missing data. This is the case for school enrollment data, where two commonly reported measures of school enrollment rates are the Net Enrollment Rate (NER) and the Gross Enrollment Ratio (GER). NER is the ratio of children of official secondary school age who are enrolled in secondary school to the population of official secondary school age children, while GER is the ratio of total enrollment in secondary school, regardless of age, to the population of official secondary school age children. NER and GER both aim to measure the same underlying quantity of school enrollment. NER can be thought of as a more refined measure of school enrollment that incorporates information about the age of children who are enrolled in school, while GER is a coarser measure of school enrollment that only requires information about the number of children enrolled. NER is more difficult to measure than GER and has a larger amount of missing data.
More generally, this type of missing data can arise if a survey is designed to collect information on a refined measure of the quantity of interest but includes the option to collect information on a coarse measure of the quantity of interest if the respondent does not provide enough information for the refined measure. For example, a survey on school enrollment might ask for information on the number and ages of children enrolled in school to calculate NER but include the option for respondents to provide only the number of children enrolled to calculate GER. In surveys that do not collect the coarse measure directly, there may exist auxiliary information from external sources, such as administrative records, that can be used as a coarse measure of the quantity of interest. This type of missing data can also occur by design in longitudinal panels to reduce respondent burden. Panel members may be asked to provide basic information for the coarse measure in every cycle of the panel, but may only be asked to provide detailed information for the refined measure in selected cycles. In all of these examples, the refined measure of the quantity of interest tends to have a greater amount of missing data than the coarse measure by nature of being more difficult to measure. If the variable of interest for subsequent analysis is the refined measure, researchers might be interested in how to best leverage information from the coarse measure to impute the variable of interest using a multiple imputation procedure.
@@
Multiple imputation, first developed by Rubin (1978, 1987), is a widely used approach for handling missing data. In multiple imputation, $M>1$ imputed values for missing observations are sampled from the posterior predictive distribution of the missing data given the observed data. The $M$ imputed values result in $M$ completed data sets, each of which consists of the observed data and one set of imputed values for the missing observations. The completed data sets can each be analyzed separately using complete data methods and the results of the analyses can be combined into one final, pooled result using combining rules from Rubin (1987) that account for both within-imputation variation and between-imputation variation. Multiple imputation approaches generally assume data is Missing At Random (MAR) as defined by Rubin (1976). For variable $\mathbf{X}$, MAR occurs when the probability of being missing depends only on the observed portion of $\mathbf{X}$. A special case of MAR occurs when data is Missing Completely At Random (MCAR) and the probability of being missing does not depend on $\mathbf{X}$. Data can also be Missing Not At Random (MNAR), which occurs when the probability of being missing depends on the missing portion of $\mathbf{X}$.
@@
We focus on the setting where the imputation of missing values and the analysis of imputed values are conducted independently. The development of multiple imputation originated in this setting, where Rubin (1977, 1978) proposed the multiple imputation framework as a way to provide imputed values for missing responses in public-use releases of large data sets from sample surveys. An appealing feature of multiple imputation for practitioners is the ability to use the same imputed data set to conduct many different analyses (Schafer (1997a); Rubin (1987)). However, using the same multiply imputed data for different analyses can lead to the analysis model being uncongenial to the imputation model in the sense of Meng (1994). An imputation model and an analysis model are congenial if there exists a Bayesian model such that (i) the posterior mean and variance from the Bayesian model for the parameter of interest are asymptotically the same as the mean and variance estimates from the analysis model in both the complete and incomplete data settings and (ii) the posterior predictive distribution of the missing data given the observed data derived from the Bayesian model is identical to the imputation model (Meng, 1994). When these two conditions are not met, the imputation and analysis models are uncongenial, and theoretical properties of the multiply imputed estimates created using the simple combining rules of Rubin (1987) may not be guaranteed (Meng (1994); Rubin (1996); Xie and Meng (2017)). Rubin’s multiply imputed variance estimator is not guaranteed to be asymptotically unbiased for the true repeated sampling variance of the multiply imputed point estimator under uncongeniality, leading to estimated confidence intervals with less than nominal coverage. Uncongeniality of the analysis and imputation models can be a regular occurrence in practice, particularly when the researchers that collect the survey data create imputed values for publication that are then used in analyses by external researchers. The external researchers may not have access to the same information or resources needed to create imputations of their own, for example if the variables used in the imputation model are not publicly available. Specifically, we consider the scenario where imputation of a variable $\mathbf{Y}$ for public release is conducted by leveraging the relationship between $\mathbf{Y}$ and an auxiliary variable $\mathbf{X}$, where $\mathbf{X}$ (and imputed values of $\mathbf{X}$) cannot be made publicly available. The multiply imputed values of $\mathbf{Y}$ are subsequently used by an external researcher in an analysis with an external outcome variable, $\mathbf{Z}$, which is unknown to the imputer and is not accounted for in the imputation process. In this scenario, the analysis model used by the external researcher is not guaranteed to be congenial to the imputation model. The ability of a multiple imputation method to perform well for uncongenial analyses is thus of interest for practitioners.
For hierarchical data, multiple imputation approaches that do not explicitly account for the hierarchical structure of the data can lead to biased results in downstream analyses (Taljaard, Donner and Klar (2008); Enders, Mistler and Keller (2016); Lüdtke, Robitzsch and Grund (2017)). Many approaches specifically for multiple imputation of hierarchical time series data have been developed (e.g. Liu, Taylor and Belin (2000); He, Yucel and Raghunathan (2011); Speidel, Drechsler and Jolani (2018); Enders, Du and Keller (2020); Grund, Lüdtke and Robitzsch (2021); among others), with two of the most widely used approaches for social science data being Amelia II and multilevel extensions of Multiple Imputation by Chained Equations (MICE). Amelia, originally developed by King et al. (2001) and extended as Amelia II by Honaker and King (2010), is a multiple imputation method designed specifically for hierarchical time series data. Amelia is based on the joint modeling approach to multiple imputation, where imputed values are sampled from a joint distribution for all variables with missing data. MICE is a multiple imputation method developed by van Buuren and Groothuis-Oudshoorn (2011) that uses the fully conditional specification (FCS) approach to multiple imputation. Rather than explicitly specifying a joint imputation model, the FCS algorithm iteratively samples from univariate conditional imputation models for all variables with missing data until convergence is reached. Several methods that account for hierarchical data structures have been implemented within the MICE framework, including the linear mixed effects method developed by Schafer and Yucel (2002).
The most commonly used methods for multiple imputation of hierarchical time series data assume that the variables in the imputation model have a linear relationship. In settings where variables have a strong nonlinear relationship and transformation to approximate linearity is not possible, these methods are misspecified. Failing to account for the nonlinear relationship can lead to imputed relationships between variables that are implausible based on substantive knowledge and biased estimates in analyses using the multiply imputed data. Nonlinear relationships in hierarchical time series data can occur regularly in the social sciences, for example if rates that aim to measure the same quantity are calculated for related populations. This is the case for school enrollment rates, where NER and GER measure school enrollment for slightly different populations of children. Another example arises for contraceptive prevalence rates, where country-level estimates of contraceptive prevalence for married and unmarried women have a nonlinear relationship when pooled across countries and times. Measurement of contraceptive use is more difficult for unmarried women than for married women, leading to a larger amount of missing data for unmarried women. A multiple imputation method for hierarchical time series data that can account for this type of nonlinear relationship could be of great utility for such applications. Existing multiple imputation methods that address more general nonlinear settings include methods based on classification and regression trees (Burgette and Reiter (2010)), functional data analysis (He, Yucel and Raghunathan (2011)), B-splines (Mbougua et al. (2013)), and generalized additive models for location, scale, and shape (de Jong, van Buuren and Spiess (2016)). We build upon this existing literature to develop a multiple imputation method for continuous hierarchical time series data that can account for a nonlinear relationship between variables that represent refined and coarse measures of the underlying quantity of interest using smoothing splines. We refer to this method as MINTS for Multiple Imputation of hierarchical Nonlinear Time Series data. We focus on the bivariate setting, where the refined measure is the variable for which imputations are desired, and the coarse measure is an auxiliary variable that is easier to measure and has a nonlinear relationship with the refined measure.
@@
This paper is structured as follows. In Section 2, we describe the motivating case study of secondary school enrollment data in further detail. Section 3 describes the proposed multiple imputation method. In Section 4, we conduct a simulation study to evaluate the out-of-sample validation performance of the proposed multiple imputation method. We compare the performance of the proposed multiple imputation method with several existing multiple imputation methods for estimation of parameters in analysis models that are uncongenial to the imputation model. In Section 5, we conduct two out-of-sample validation exercises using the motivating data set on secondary school enrollment rates, where we evaluate predictive performance for estimation of individual missing values and estimation of parameters in uncongenial analysis models. We also create multiple imputations for the full school enrollment data set and make available 40 multiple imputations for NER created using MINTS. Section 6 includes further discussion and comparisons of the proposed multiple imputation method with existing methods. Finally, we summarize the findings of this paper in Section 7.
@@ 2 Motivating Case Study: Secondary School Enrollment Rates
The UNESCO Institute for Statistics collects internationally comparable data on education indicators on an annual basis for all countries of the world, based largely on survey and administrative data (UNESCO Institute for Statistics (2023)). The World Bank combines this education data with population data from the United Nations Population Division to create estimates of two types of enrollment rates: the Net Enrollment Rate (NER) and the Gross Enrollment Ratio (GER) (World Bank (2021)). We focus on secondary school enrollment. NER is the ratio of children of official secondary school age who are enrolled in secondary school to the population of official secondary school age children. NER is bounded between 0% and 100% and both the numerator and denominator reflect children of official secondary school age. GER is the ratio of total enrollment in secondary school, regardless of age, to the population of official secondary school age children. The numerator and denominator for GER potentially represent different populations, where children who are not of official secondary school age can be counted in the numerator but not in the denominator. Thus, GER can be greater than 100% if children who are enrolled in school are not of official school age. The two measures of enrollment have a strong nonlinear relationship and are subject to the boundary NER $\leq$ GER.
For substantive analyses, one measure of enrollment may be preferred over the other. NER can be thought of as a demographic rate, where the numerator counts the number of enrollments for the population of children of official secondary school age in a given year and the denominator counts the person-years lived in that population for the given year. NER can thus be preferable over GER for demographic analyses. However, historical time series of NER tend to have more missing values than time series of GER, as measurement of NER is more difficult than measurement of GER due to requiring knowledge of the age distribution for children enrolled in school. Measurement of GER is comparatively easy, as GER can be calculated using only the number of children who are currently enrolled. For school systems that do not have robust recordkeeping systems and countries that do not have good vital registration systems, knowledge of the age of all enrolled children can be difficult to obtain. The greater availability of estimates of GER and the strong relationship between NER and GER motivates the desire to impute missing values of NER using the relationship between NER and GER.
@@
Figure 1 shows a scatter plot of the complete cases for NER and GER. The superimposed line illustrates the B-spline of degree 1 fit to the complete cases using the A-splines methodology described in Section 3.2.5. There is a nonlinear relationship between NER and GER, with a shift in trend occurring around GER $=100$. The variation of NER about the fitted spline also appears to vary with GER, with smaller variability at the lowest levels of GER and larger variability around GER $=100$.
Figure 1: Scatter plot of complete cases for NER and GER from secondary enrollment data set superimposed with the B-spline of degree 1 fit using A-splines.
Time series of NER and GER are shown in Figure 2 for Afghanistan, Belgium, Spain, and Nigeria. Afghanistan is an example of the most common type of pattern seen for individual countries, where there is a larger number of observed values for GER compared to NER. Only one country, Brazil, has the opposite pattern with one more observed value of GER than observed values of NER. Belgium and Spain are two examples of countries where the nonlinear relationship between NER and GER is visible in the time series for the individual country. Both Belgium and Spain have relatively few missing values for GER, but have large stretches of time with no observations for NER. Finally, Nigeria is an example of a country that has some observed values for GER but has no observed values for NER. There are 14 countries in the school enrollment data set that, like Nigeria, have at least one observation for GER but no observations for NER.
Figure 2: Observed values of NER and GER for selected countries from secondary enrollment data set. Open and solid circles indicate observed values for NER and GER, respectively.
@@ 3.1 Notation
Let the matrices $\mathbf{R}^{X}$ and $\mathbf{R}^{Y}$ denote the response matrices for $\mathbf{X}$ and $\mathbf{Y}$, respectively. If an element $(c,t)$ is observed, then the corresponding element of the response matrix equals 1. For example, if $X_{c,t}$ is observed, then $R_{c,t}^{X}=1$. If $X_{c,t}$ is missing, then $R_{c,t}^{X}=0$. The matrices $\mathbf{X}$ and $\mathbf{Y}$ can also be written in terms of their observed and missing portions. For example, $\mathbf{X}=[\mathbf{X}_{mis},\mathbf{X}_{obs}]$ where the observed portion of $\mathbf{X}$ is denoted $\mathbf{X}_{obs}$ and the unobserved portion of $\mathbf{X}$ is denoted $\mathbf{X}_{mis}$.
For the school enrollment data, the auxiliary variable $\mathbf{X}$ is GER and the variable of interest $\mathbf{Y}$ is NER. The enrollment data includes a total of $C=202$ countries, where the assignment of countries to indices $c\in 1,...,C$ was done alphabetically by country name. There are $T=51$ years in the enrollment data set, where $t=1$ corresponds to the year 1970 and $t=51$ corresponds to the year 2020.